government agency has a construct 30 kilometre of air strip in 6 month it takes 450 workers to do the job but the order is extended by another 10 kilometre the agency must also finish the task in 4 months how many more workers put the agency hire to complete the strip on time
1
Expert's answer
2019-12-09T12:15:42-0500


"...30 kilometers (km) of air strip in 6 month it takes 450 workers to do the job"

= > 5 km of air strip in 1 month it takes 450 workers to do the job

Still left 30+10 = 40 (extended road)

40 km for 4 months.

10 km for 1 months

"5 km of air strip in 1 month it takes 450 workers to do the job "

=> for 2*5 km of air strip in 1 month it takes 2*450 workers to do the job.

The agency must have

2*450 workers = 900 workers.

To complete the strip on time the agency should hire

900-450 = 450 workers.